• The ADSC Q315r was delivered on Nov 6 and was installed at BL-5A. We started the operation on Nov 8.(2008/11/10)
  • The ADSC found trouble on Q315r during their work, so that the delivary is delayed more. After Oct 28, the BL-5A will be operataed with Quantum 4R instead. Sorry for the inconvenience.(2008/10/27)
  • Due to the delay in upgrading and installing the the ADSC Quantum 315r, the detector of BL-5A is now ADSC Quantum 210r. Sorry for the inconvenience. It will be returned back on Oct 27.(2008/10/14)
  • The ADSC Quantum 315 of BL-5A was shipped because of the upgrade and overhaul. It will be returned back as the Q315r this autumn.(2008/07/01)
  • The table of the "Comparison of specifications" was updated.(2008/02/04)
  • We installed a new detector, ADSC Q270, at the BL-17A and executed the first test experiment. We will start test operation tomorrow.(2007/06/12)
  • The beamtime of 26th and 27th of May were canceled due to PF ring trouble. Both PF and PF-AR were recovered and are now operational.(2007/05/30)
  • The beamtime before 23rd of May were canceled due to PF-AR ring trouble. Now we are recovering the trouble and will start user operation after 23rd.(2007/05/16)
  • The BL-17A photo page, "Current photographs of the BL-17A", was updated. Let's have a look of it on the BL-17A page.(2007/04/18)
めもCGI by Chama-Net